The width of a rectangle is 23 cm. Find all possible values for the length of the rectangle if the perimeter is at least 704 cm.

Answer by stanbon(75887)

The width of a rectangle is 23 cm. Find all possible values for the length of the rectangle if the perimeter is at least 704 cm.
-----
P = 2(L+W)
----
Inequality::
2(L+23) >= 704
----
L+23 >= 352
====
Length >= 329 cm
------------------------
Cheers,
Stan H.